OpenStudy (anonymous):

In a certain candy store, 3 pounds of candy and 2 pounds of mints cost $10.80, and 1 pound of candy and 3 pounds of mints cost $5.35. What is the cost per pound of the mints? mints are 65¢ per lb mints are 75¢ per lb mints are 85¢ per lb

OpenStudy (anonymous):

Let c and m be the cost, $ per lb, of candy and mints respectively. Solve the following equations for c and m. {3 c + 2 m = 10.80, 1 c + 3 m = 5.35} {c = $3 .10/lb, m = $0 .75/lb} mints are 75¢ per lb